# Understanding and Mitigating Compositional Issues in Text-to-Image Generative Models

### v1

@@ Acknowledgement
Appendix A Limitations
In this paper, we have thoroughly analyzed one of the key reasons why Stable Diffusion struggles to generate compositional prompts and proposed a lightweight method to mitigate this issue. However, there remains significant room for improvement in this area. Our approach focuses on improving the text encoder, which we identified as a major source of error. There are potentially other sources of the issue within the entire generative model pipeline that need to be explored. Additionally, our method involves a small fine-tuning step using a simple linear projection. Future work could explore alternative approaches, such as more sophisticated fine-tuning techniques, advanced attention mechanisms, or hybrid models that integrate multiple strategies.
Appendix B Optimizing the Text-embedding of a Subset of Tokens
@@
Appendix C Source (i) : Erroneous Attention Contributions in CLIP
C.1 Attention Contribution
@@
C.2 Zero-shot Attention Reweighting
@@
Appendix D Training setup
Appendix E CLP and WiCLP Visualization
In this section, we provide additional visualizations comparing CLP, WiCLP, and baseline models in Figures 12, 13.
Appendix F Visualization of Cross-Attentions
In this section, we provide additional cross-attention map visualizations in Figures 12 and 13.
Appendix G Visualization of Switch-Off
In this section, we present more qualitative samples illustrating the effect of Switch-Off at different timestep thresholds for various prompts in Figures 14 and 15.
@@ Appendix G Visualization of Switch-Off
Figure 10: Caption
@@
Figure 11: Caption
@@
Appendix H Choice of Window Length in WiCLP
One might suggest that instead of using token-wise linear projection (CLP) or a window-based linear projection with a limited window (WiCLP), employing a linear projection that considers all tokens when finding a better embedding for each token might yield better results. However, our thorough quantitative study and experiments tested various window sizes for WiCLP. We found that using a window size of 5 achieves the highest performance.

### v2

@@ Acknowledgement
Appendix A Optimizing the Text-embeddings of a Subset of Tokens
@@
Appendix B Source (i) : Erroneous Attention Contributions
B.1 Attention Contribution
@@
Figure 13: Overview of different architectural designs for the projection layer. Each design varies in structure and functionality, influencing how input features are transformed and integrated into the model.
Figure 14: Comparison of performance, speed, and parameter efficiency across different projection architectures. Each projection is represented by a different color, with marker size indicating its relative number of parameters. The x-axis represents the time spent during the backward pass, while the y-axis shows the average performance across the three attribute binding categories: color, texture, and shape.
B.2 Zero-shot Attention Reweighting
Inspired by attention mechanism shortcomings of CLIP text-encoder, we aim to improve compositionality of CLIP-based diffusion models by zero-shot reweighting of the attention maps. Specifically, we apply a hand-crafted zero-shot manipulation of the attention maps in certain layers of the CLIP text-encoder to effectively reduce unintended attentions while enhancing meaningful ones. This zero-shot reweighting is applied to the logits before the Softmax layer in the last three layers of the text-encoder. More precisely, we compute a matrix $M\in\mathbb{R}^{n\times n}$ and add it to the attention logits. For each head $h$, the new attention values are computed and then propagated through the subsequent layers of the text encoder:
We set the values in $M$ by considering the ideal case where no incorrect attentions occur in the mechanism. For example, for prompt “a green bench and a red car", we ensure that the token "car" does not attend to the token "green" by assigning a sufficiently large negative value to the corresponding entry in matrix $M$.
@@ B.2 Zero-shot Attention Reweighting
Applying zero-shot attention reweighting with matrix $M$ on $780$ compositional prompts of the color category of T2I-CompBench dataset, we achieved a $2.93\%$ improvement in VQA scores.
B.3 Experiments with LLaMa3 8B
LLaMa3 CLIP
color last layer 0.015 0.657
all layers 0.081 0.187
texture last layer 0.033 0.696
all layers 0.066 0.213
[t]
Table 2: Unintended attention rate in LLaMa3 8B vs CLIP. LLaMa3 shows significant less unintended attentions.
FID Score
SD v1.4 24.33
SD v1.4 + WiCLP 25.40
SD v2 23.27
SD v2 + WiCLP 27.40
GORS 30.54
Table 3: Comparison of FID scores between the baseline models and WiCLP using Switch-Off with $\tau=800$, as well as the GORS approach.
We explored the analysis of attention contributions to identify unintended attention in LLaMa3 8B, which utilizes a more advanced text encoder specifically designed for language modeling and pretrained on large-scale text corpora. Table 2 reports the rate of unintended attention across prompts in the color and texture datasets. The results demonstrate that unintended attention occurs less frequently in more advanced text encoders, further emphasizing the limitations of the CLIP text encoder.
B.4 Models with T5 text-encoder
We conducted experiments to measure the VQA score on the color dataset for models that use T5 as their text encoder. DeepFloyd achieved a score of $0.604$, which is significantly higher than that of SD-v1.4. Additionally, DeepFloyd-I-M, which employs a smaller first-stage UNet compared to DeepFloyd, obtained a score of $0.436$, also surpassing the SD-v1.4 score.
Appendix C Ablation on the Projection Layer Architecture
In this section, we discuss the architectural choices we considered for the projection layer. As illustrated in Fig. 13, we experimented with five different architectures, ranging from simple linear networks to more advanced attention-based models.
The simplest architecture, shown in Fig. 13 (a), consists of a single linear or windowed linear convolutional layer that transforms the input into a new embedding. Building on this, we introduced a residual connection from the input to the output, resulting in the architectures depicted in Fig. 13 (b) and (c). These correspond to CLP and WiCLP, which we discuss in detail in Section 4. Additionally, we explored the effect of incorporating non-linearity into these projection layers, as shown in Fig. 13 (d).
To design a more sophisticated architecture, we developed an attention-based projection model, which we refer to as the Shallow Transformer Projection (STP). STP consists of stacked attention blocks similar to those in transformer architectures but with a key difference: we introduce a Token-wise Scalar Network that dynamically controls the extent to which each token’s representation is influenced by the projection block (see Fig. 13 (e)).
For a comprehensive comparison, we evaluated these projection models across various diffusion model variants, ranging from SD v1.4 and SD v2 to more recent models like SDXL. While all projection methods improved upon the baseline, CLP and WiCLP demonstrated the most noticeable gains despite their parameter efficiency. This highlights that CLP and WiCLP not only offer a lightweight and computationally efficient solution but also achieve superior compositional improvements.
To ensure robust and generalizable quantitative results, we evaluated different projection architectures across all attribute binding categories of T2I-CompBench—color, texture, and shape—on SD v2. The results, presented in Figure 10, demonstrate that CLP and WiCLP achieve significantly higher performance gains while maintaining parameter and speed efficiency. Additionally, Figure 14 compares the projections in terms of performance, inference speed, and parameter count. Notably, increasing model complexity, such as with the Shallow Transformer projection, does not necessarily lead to better performance. In contrast, CLP and WiCLP strike an optimal balance, offering superior compositional attribute binding while remaining highly efficient in both speed and parameter usage.
Appendix D Experiments
D.1 Training setup
D.2 Extended Evaluation
Human Evaluation We conducted a human evaluation in which participants compared images generated by SD v1.4 and SD v1.4 + WiCLP, selecting the image that best matched the given prompt (Figure 22). Five evaluators were presented with 200 randomly selected image pairs, evaluating a total of $1000$ image-caption pairs.
TIFA Metric. To provide a more comprehensive evaluation, in addition to the disentangled BLIP-VQA score proposed by (Huang et al., 2023), we also incorporate the TIFA metric (Hu et al., 2023). TIFA (Text-to-Image Faithfulness Evaluation with Question Answering) is an automated evaluation method that measures how faithfully a generated image corresponds to its textual input via visual question answering (VQA). It generates multiple question-answer pairs from the text input using a language model, then evaluates image faithfulness by determining whether existing VQA models can accurately answer these questions based on the image. As a reference-free metric, TIFA offers fine-grained and interpretable assessments of image quality.
Using TIFA, we observed that SD v1.4 and SD v2 achieved scores of $0.6598$ and $0.7735$, respectively. Notably, the scores for WiCLP applied on top of SD v1.4 and SD v2 improved to $0.7462$ and $0.8133$, respectively, demonstrating the enhanced performance of our approach.
FID Score Comparison Our method results in a modest increase in FID score on MS-COCO prompts compared to the base models, as shown in Table 3. However, this increase is less pronounced than in other baselines—for example, SD v2 + WiCLP scores 27.40, whereas GORS reaches 30.54.
D.3 Robustness and Generalizability of WiCLP
To validate the robustness and generalizability of our method, we trained WiCLP across all categories simultaneously using the same setup as when training WiCLP on individual categories separately. For SD v2, this resulted in VQA scores of 0.6311 for color, 0.5728 for texture, and 0.4620 for shape—a slight ($\scriptstyle\sim 2\%$) decrease compared to training on individual categories. Despite this minor drop, the model still significantly outperforms the baseline, demonstrating its strong generalizability and robustness.
Additionally, to assess the generalizability of WiCLP to other text encoders, we conducted experiments on PixArt-$\alpha$ and DeepFloyd, both of which use a T5 text encoder. The results, presented in Table 1 and Appendix D.4, demonstrate that WiCLP is highly effective for these models, further highlighting its robustness and broad applicability.
D.4 Generalization of WiCLP to T5-Based Encoders
To assess WiCLP’s applicability to text encoders beyond CLIP, we tested it on DeepFloyd-I-M and PixArt-$\alpha$, both of which use T5-XXL as their text encoder. Our results confirm that the linear head tuning approach remains effective, yielding significant improvements across different architectures.
To establish that these models have room for improvement, we conducted interpretability experiments as outlined in Section 3.2. For instance, with the prompt “a red book and a yellow vase”, the baseline DeepFloyd-I-M model achieved a BLIP-VQA score of 0.4350. By optimizing the text embedding space with our method, this score significantly increased to 0.9121, demonstrating that the T5 text encoder’s output space can be further refined for better compositional generation in these models.
Table 1 presents the results of applying WiCLP to PixArt-$\alpha$, showing noticeable improvements over the baseline. Similarly, applying WiCLP to DeepFloyd-I-M enhanced BLIP-VQA performance across all 300 evaluation prompts, increasing the score from 0.4636 to 0.5155. These results further reinforce WiCLP’s generalizability and effectiveness across diverse text encoders.
D.5 CLP and WiCLP Visualization
In this section, we provide additional visualizations comparing CLP, WiCLP, and baseline models in Figures 18, 19.
D.6 Visualization of Cross-Attentions
In this section, we provide additional cross-attention map visualizations in Figures 18 and 19.
@@ D.6 Visualization of Cross-Attentions
D.7 Visualization of Switch-Off
In this section, we present more qualitative samples illustrating the effect of Switch-Off at different timestep thresholds for various prompts in Figures 20 and 21.
D.8 Choice of Window Length in WiCLP
One might suggest that instead of using token-wise linear projection (CLP) or a window-based linear projection with a limited window (WiCLP), employing a linear projection that considers all tokens when finding a better embedding for each token might yield better results. However, our thorough quantitative study and experiments tested various window sizes for WiCLP. We found that using a window size of 5 ($s=2$) achieves the highest performance.
@@ D.8 Choice of Window Length in WiCLP
Figure 16: Qualitative comparison between the baseline and our projection methods (CLP and WiCLP).
@@
Figure 17: Qualitative comparison between the baseline and our projection methods (CLP and WiCLP).
@@
Figure 22: A sample from the human evaluation study, where participants were presented with a pair of images and a caption. They were asked to select the image that best represented the caption or choose ’both’ if the images equally captured the caption’s meaning.